OpenStudy (anonymous):

Derek distributed 30 pieces of apple to his friends. If each apple slice was 1/8 of an apple, how many apples did he distribute to his friends?

jimthompson5910 (jim_thompson5910):

There are 30 pieces, all 1/8th of an apple, so you have 30*(1/8) = 30/8 = (24+6)/8 = 24/8 + 6/8 = 3+3/4 = 3 3/4 So he distributed 3 3/4 apples to his friends Basically he gave out 3 whole apples and 3/4 of an apple as well